OpenStudy (anonymous):

Using point slope form y-y1 = m(x-x1) write an equation for the line: a. With slope -2, containing point (3,4) c. With slope 0, containing the point (4,2)

OpenStudy (triciaal):

a. m = -2 point (3, 4) means x1 =3 and y1 = 4 substitute in y-y1 = m(x-x1) b. does not exist c. follow the same format as a.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ 4}})\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 2}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= -2\\ slope = {\color{green}{ m}}= 0\\ \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)